Determine the refractive index for a material when the speed of light in that material is 35000000 m/s. Answer to 1 decimal places.
1
Expert's answer
2020-06-22T11:10:14-0400

Given, that speed of light in a certain medium is "v=3.5\times10^7m/s"

Since, refractive Index is defined by

"\mu=\frac{c}{v}\implies \mu=\dfrac{3\times 10^8}{3.5\times 10^7}\approx8.6"